hello solvers ..i need some help from yours..hope u will answer it now, thanks..-1/2(x-6)+1/3(x+5)=x+8 i been solving it ten times but i couldn't get the rigth answer.. i need some help..thanks.
Answer by stanbon(75887)

Multiply thru by "6" to get rid of the fractions; then
-3(x-6)+2(x+5)=6(x+8)
-7x = 20
x = -20/7
Cheers,
Stan H.